# REASSEMBLE: A Multimodal Dataset for Contact-rich Robotic Assembly and Disassembly

### v1

@@ Appendix A Data visualization tool
Figure 10 showcases the data visualization tool developed using the ReRun Python package. The top-left widget displays all action segments and success annotations, with the current action highlighted. The center and right widgets in the top row show one of the external and the wrist camera perspectives, respectively. In the second row, the first widget presents the second external perspective, the center widget displays the event camera mono image, and the rightmost widget visualizes the event data. The third row illustrates the waveforms from all microphones. The remaining widgets depict the robot’s proprioceptive data, including compensated forces and torques, joint and gripper states, as well as the end-effector pose and orientation.
Figure 10: Snapshot from the dataset visualization tool.

### v2

@@ Appendix A Data visualization tool
Figure 12 showcases the data visualization tool. Green-marked widgets display all action segments and their success annotations, with the current action underlined. Red widgets show external and wrist camera views, while the bottom-right widget visualizes event data. Blue widgets illustrate audio waveforms from all microphones, and purple ones present proprioceptive data, including compensated forces and torques, joint and gripper states, and end-effector pose and orientation. The orange-marked timeline at the bottom displays the exact timestamps for each sensory modality. The rightmost widget shows the layout structure (Blueprint) of the visualization tool. The tool is accessible online via the project website.
Figure 12: Snapshot from the dataset visualization tool. Best viewed in color.
Appendix B Dataset structure
Figure 11 shows the structure of a single demonstration file in the REASSEMBLE dataset. Each demonstration is stored in an individual file, identified by the date and time the demonstration began. The HDF5 (.h5) format distinguishes between two types of structures: datasets, which contain the raw data, and groups, which are folder-like elements that can contain multiple datasets or subgroups. In Figure 11, groups are represented by folder icons, and datasets by file icons. Video, audio, and event data are stored directly within the main group of the file. Video and audio are stored as encoded byte strings to reduce memory usage, while event data is stored as arrays. The robot’s proprioceptive data is stored in the robot_state group as arrays. Since each sensor modality is recorded at its own measurement frequency, the arrays have different lengths. To synchronize the sensory information, each modality’s timestamps are stored in the timestamps group. Action segment information is stored in the segments_info group. Each segment has its own subgroup, identified by its order in the demonstration. Each segment contains a starting timestamp, ending timestamp, a success flag, and a natural language description of the action. Skills are stored within a low_level subgroup inside each segment, and follow the same structure as the high-level annotations. Additionally, each HDF5 file has an associated JSON file containing the poses (positions and quaternions) of the cameras and the board. Code for loading and saving the HDF5 files is available on GitHub.
(a) BNC connector and socket. Approx. diameter: 15 mm.
(b) D-SUB connector and socket. Approx. size: 60$\times$20 mm.
(c) Waterproof connector and socket. Approx. size: 50$\times$35 mm.
(d) Gears: small, medium, and large (from left to right), mounted on a plate. Approx. diameters: 20 mm, 40 mm, and 60 mm.
(e) Nut 4 and its socket. Approx. diameter: 25 mm.
(f) Top: round pegs. Bottom: square pegs. Round peg diameters (sizes 1–4): 4 mm, 8 mm, 12 mm, 16 mm. Square peg dimensions (sizes 1–4): 4 mm, 8$\times$7 mm, 12$\times$8 mm, 16$\times$10 mm.
(g) USB connector and socket. The connector is held in a custom holder.
(h) Ethernet connector and socket. The connector is held in a custom holder.
Figure 13: Objects and connectors used in the REASSEMBLE dataset.
Appendix C Objects
Figure 13 shows photos of the objects, connectors, and holders used in the REASSEMBLE dataset. We use 17 objects from the NIST task board 1 in the REASSEMBLE dataset:
  • 5 electrical connctors: Ethernet, USB, waterproof, BNC, D-SUB
  • 3 gears: small, mediun, and large
  • 1 M16 Nut
  • 4 round pegs: 4mm, 8mm, 12mm, and 16mm
  • 4 square pegs: 4mm x 4mm, 8x7, 12x8, 16x10